OpenStudy (maheshmeghwal9):

Let A,B be the roots of the equation x^2-px+r=0 & A/2,2B be the roots of this equation x^2-qx+r=0. Then what is the value of {r} ?

OpenStudy (anonymous):

in terms of????

OpenStudy (anonymous):

In terms of p and q????

OpenStudy (asnaseer):

if a quadratic equation has roots \(r_1, r_2\) then it can be written as:\[(a-r_1)(x-r_2)=0\]use this to find your solution.

OpenStudy (maheshmeghwal9):

p & q.

OpenStudy (asnaseer):

sorry, first bracket should be \(x-r_1)\)

OpenStudy (anonymous):

(5pq-q^2-4p^2)/9

OpenStudy (anonymous):

Use this property to solve this sum Sum of roots of quadratic equation \[ax^2+bx+c=0\] is -b/a

OpenStudy (anonymous):

am i righr????

OpenStudy (maheshmeghwal9):

but how?

OpenStudy (anonymous):

AB=r A+B=p A+4B=2q solving above 2 we get B=(2q-p)/3 A=4p-2q)/3 put value of A and B in top equation we get r=(10pq-4q^2-4p^2)/9

OpenStudy (maheshmeghwal9):

but ans. is \[2/9(2q-p)(2p-q).\]

OpenStudy (anonymous):

I have writen this only if u open the brackets

OpenStudy (maheshmeghwal9):

and which top eq. are u talking about?

OpenStudy (anonymous):

AB=r

OpenStudy (maheshmeghwal9):

Oh ,I See Thanx a lot:)

OpenStudy (anonymous):

it was pleasure nice question